# Revisiting In-context Learning Inference Circuit in Large Language Models

### v3

@@ 4 Induction Circuits in Large Language Models
This section mainly shows how LMs utilize the encoded linear text representations in induction circuits with a typical 2-step form (Singh et al., 2024b): Forerunner Token Heads merge the demonstration text representations in the forerunner token into their corresponding label tokens with a selectivity regarding the compatibility of demonstrations and label semantics. Induction Heads copy the information in the label representations similar to the query representations back to the query. Such operations are conducted on task-specific subspaces, enabling LMs to solve multiple tasks by multiplexing hidden spaces.
@@ 4.1 Step 2, Forerunner Token Head: Copy from Text Feature to Label Token
This subsection mainly examines and measures the forerunner token heads, which copy the information in the forerunner into label tokens. We investigate the interaction between the forerunner tokens and label tokens, and focus on how the representations are merged, especially when the semantics of labels and text are disjoint, towards an explanation of why ICL are robust to wrong labels.
@@
Text Representations are Copied without Selectivity. For each attention head, we extract the attention score [^7] This paper use notation $\alpha_{k\rightarrow q}$ to denote attention score with attention query $q$ and attention key $k$. $\alpha_{s_{t}\rightarrow y_{t}}$ from a forerunner token $[s_{t}]$ (as attention key) to the corresponding label token $[y_{t}]$ (as attention query). We then mark the head with $\alpha_{s_{t}\rightarrow y_{t}}\geqslant 5/n_{t}$ ($n_{t}$: the length of tokens before $[y_{t}]$) as a Forerunner Token Head and count them in each layer. The results are shown in Fig. 5 (Middle, “Correct Label”), where the peak matches the copy period in Fig. 5 (Left). Moreover, to investigate the influence of the correctness of label tokens, we replace $[y_{t}]$ with a wrong label token [^8] For example, for a label space of “positive” and “negative”, if $[y_{t}]$ is “positive”, we replace it to “negative”., where the results in Fig. 5 (Middle, “Wrong Label”) are almost identical to the correct-label setting, suggesting that the forerunner token heads don’t show selectivity toward the semantic consistency between input text and labels simply merge the input text representations into the label tokens. Furthermore, we find (Appendix D) that the copy processing is inherent: during the copy processing, LMs establish strong attention connections to preceding tokens, regardless of whether these copied tokens are forerunner tokens, which indicates that such copy processing is a universal inference behavior developed during pre-training, rather than being triggered by the special tokens in the in-context learning (ICL) input, while still aiding ICL processing (discussed in §5.1).
@@
Hidden States of Label Tokens are Joint Representations of Text Encodings and Label Semantics. Given the findings above, we probe the content of hidden states of label tokens $[y_{t}]$, i.e., how the copied text representation interacts with the original label semantics. We first train two centroid classifiers to predict the corresponding label $y_{t}$: (1) $\mathcal{C}_{f}$ trained on the hidden states of forerunner tokens $[s_{t}]$ and (2) $\mathcal{C}_{l}$ trained on the hidden states of label tokens $[y_{t}]$. To check whether the label tokens include the information of forerunner tokens, we use $\mathcal{C}_{f}$ to predict the label on the hidden state of label token $[y_{t}]$ in Fig. 5 (Right, solid). It shows that, during the copy processing, high classification accuracies can be achieved both on the correct label tokens and wrong label tokens, suggesting that the text features in the forerunner tokens can be partly and linearly detected in the label tokens. Moreover, results using $\mathcal{C}_{l}$ (dotted line) shows extreme results, suggesting the label information remains in the label token. So, we can conclude that: hidden states of label tokens are joint representations of label semantics and text representations. Moreover, interestingly, the accuracies from $\mathcal{C}_{f}$ shown in Fig. 5 (Right) decline after layer 35, suggesting that the information sharing between the forerunner tokens and the label tokens ends in later layers, which aligns with the results in Fig. 5 (Middle).
@@
Label Denoising is Conducted on the Overlap of Label Semantics and Text Representations. Notice that in Fig. 5 (Right, solid), compared to the typical results predicted on the forerunner tokens, accuracies are improved on the correct label and suppressed on the wrong label, which suggests that information consistent with the label semantics is easier to be enhanced by the label tokens and vice versa, showing a feature selectivity on the consistency between the text representations and the label semantics. Given the observation that the information on label semantics and text features can be extracted separately and linearly, we can confirm that these two kinds of information are located in different sub-spaces of the hidden states, and linearly merged by the attention operation of forerunner heads. Moreover, given the fact that there is no selectivity is observed in the copy behavior of the forerunner token head (Fig. 5 (Middle)), it is intuitive that the feature selectivity shown in Fig. 5 (Right) comes from the arithmetical interaction of feature vectors on the overlap of sub-spaces between the label semantics and text features, making ICL stable against label noise (Min et al., 2022). Moreover, as mentioned by Wei et al. (2023), large models show poorer stability against label noise, and we infer that the larger hidden dimensions in larger models lower the overlap between the sub-spaces of label semantics and text representations to reduce the interaction.
@@
Figure 5: Hidden states copy magnitude from forerunner tokens to label tokens against layers. Left: Kernel alignment between the forerunner token (the copy source) and the label token of the next layer (the copy target). Middle: Curves: The count of marked forerunner token heads with correct and wrong labels; Colored Areas: The maximum attention scores from forerunner token to query (copy magnitude) with correct and wrong labels (detailed attention head statistical data is in Appendix H.1). Right: Centroid classifier results predicted on the hidden states of correct and wrong label tokens, on SST-2 and MR. (Solid: Predicted by classifiers $\mathcal{C}_{f}$ trained on hidden states of forerunners. Dotted: Predicted by classifiers $\mathcal{C}_{l}$ trained on hidden states of label tokens.)
@@ 4.2 Step 3, Induction Head: Feature Retrieval on Task Subspace
This subsection examines the existence of the aforementioned induction heads, which retrieve similar label token features as the queries’ forerunner feature, and copy the retrieved features back to the query. We claim the necessity of multi-head attention in this process: correct feature retrieval can only be conducted on the subspace of the hidden space, which is captured by some attention heads.
@@
Induction is Correct in Minority Subspaces. Similar to Fig. 5 (Middle), we mark (1) the attention heads with the sum of attention scores from all the label tokens $[y_{1}],\dots,[y_{k}]$ (as attention keys) to the query’s forerunner token $s_{q}$ (as attention query) in the demonstration more than $5k/n_{t}$ as induction heads, and (2) attention heads with the sum of scores from all the correct label tokens more than $5k/|\mathbb{Y}|n_{t}$ as correct induction heads ($\mathbb{Y}$ is a label space). We show the number of both kinds of induction heads in Fig. 6 (Left, detailed head statistics in Appendix H.1), where a unimodal pattern is observed later than the copy processing of Step 2. Moreover, more than half of the induction heads are not correct ones, suggesting that task-specific feature similarity can only be caught on some induction subspaces (defined by low-rank transition matrix $W_{Q}^{h\top}W_{K}^{h}$ of correct induction head $h$). We enhance this claim in Fig. 6 (Middle) (details in Appendix A.4), where both vanilla attention (without transformation and head split) and attention scores averaged among heads show low assignment on correct label tokens, while some heads show considerable correctness. Considering the average value, the majority of attention heads almost randomly copy label token information to the query, causing the prediction biased to the frequency of labels in the prompt (Zhao et al., 2021). As the reason, we infer that the hidden states are sufficient (Fig. 3) but not minimum for ICL, where redundant information interferes with the similarity calculation of attention.
@@
Figure 6: Measurements for induction heads. Left: The count of marked induction heads and correct induction heads against layers. Middle: The correctness of attention assignment (the sum of assignment towards correct label tokens normalized by assignment towards all the label tokens). (Vanilla Attention: attention scores directly calculated on full dimensionality. Head Average: the averaged attention scores among all the heads. Best Ind. Head: the scores of attention head with the most correctness.) Right: The correct induction heads overlap of all dataset pairs.
@@
Figure 7: Label representations of $k$ demonstrations visualized on Left 4: correct and Right 4: wrong induction head, on one sample of SST-2 (see Appendix H.3). $\circ$: “positive” label, $\times$: “negative” label, $\blacktriangle$: zero vector. Color: attention assigned to query, negative to positive (cartography: Appendix A.5).
@@
Some Induction Subspaces are Task-specific. We check if different tasks share the same induction subspaces based on the overlap of the correct induction heads across different datasets. Given $n_{\mathcal{D}}(h)$, the number of times $h$ is marked as correct induction head on dataset $\mathcal{D}$, the overlap rate $S$ is defined as:
$S(\mathcal{D}_{1},\mathcal{D}_{2})=\frac{2\sum_{\forall h}\mathrm{min}[n_{\mathcal{D}_{1}}(h),n_{\mathcal{D}_{2}}(h)]}{\sum_{\forall h}n_{\mathcal{D}_{1}}(h)+n_{\mathcal{D}_{2}}(h)}.$ (1)
@@
The results are shown in Fig. 6 (Right), where: (1) A significant overlap of induction heads indicates that a part of correct induction heads is inherent in the model, built by the pre-training process (Reddy, 2024; Singh et al., 2024b). (2) Such overlap is not fully observed, suggesting that some induction subspaces are task-specific: input texts evoke task-specific attention in induction heads, enabling the anisotropy multiplex of different subspaces in the hidden spaces to transmit relevant information for various tasks. Therefore, we can restore ICL to implicit end-to-end multi-task learning with hidden state multiplexing since they also use various task heads on common bottom network layers and informative hidden states (Zhang & Yang, 2021).
@@
Demonstrations Saturate on Induction Subspace. We visualize the demonstrations’ label token representations mapped on the induction subspaces by the transition matrix $W_{Q}^{h\top}W_{K}^{h}$ and principal component analysis in Fig. 7, indicating: (1) Compared to a correct induction head, a wrong induction head is easier to map label representations linearly inseparably. (2) In the early stage of demonstration ($k=1\rightarrow 2$), when a new demonstration is given, the morphology of attention assignment towards query changes significantly (shown as background color in Fig. 7; see §A.5 for details), while in the late stage ($k=15\rightarrow 16$), attention assignment morphology is stable. This can explain the demonstration saturation (Agarwal et al., 2024; Bertsch et al., 2024): the performance is submodular against the demonstrations. Intuitively, since demonstrations follow a prior distribution, representation of a new demonstration is likely to be located within the closure of existing demonstrations, making it less contributable to the attention assignment in the induction subspace.
@@ 5 Label $y_{i}$rightarrow Query Forerunner $s_{q}$
To demonstrate that our 3-phase circuit dominates or at least participates in ICL process, we disconnect the related attention connection of each step in the proposed circuit (see Appendix A.7 for details), and test the accuracies without such connections as shown in Table 1. The results show that: compared to the controlled results where trivial connections are ablated, when the non-trivial connections designated by the proposed circuit are ablated, accuracies of ICL significantly decrease, supporting the existence of our circuit. However, the result doesn’t fully match expectations, for example, the result without induction (line 5) should be consistent with zero-shot (line 6), since all the expected communication from demonstration to query is intercepted, but that’s not the case; and the contribution of Step 1 in later layers are unexpectedly high, indicating the existence of some bypass mechanisms parallelly contributing to ICL accuracies.
@@ 5.2 Bypass Mechanism
Parallel Circuits. Multiple 3-step circuits can execute in parallel, that is, one layer can assign multiple inference functions to different heads, causing dispersion and deserialization as shown in Fig. 8, where a narrow (fewer heads) and deep (more layers) model is more likely to generate localized inference, and vice versa.
Direct Decoding. Residual connection to output embedding allows intermediate hidden states to be decoded directly. Intuitively, a shortcut from Step 1 encodings to the LM head enables ICL with zero-shot capacities, since we have confirmed that encoded representations are informative for ICL tasks (§3.2), while the decoding methods should be selected carefully (Cho et al., 2024) (e.g. with essential calibration). On the other hand, shortcuts from insufficiently encoded features may lead to meaningless information decoded by language model heads, causing prediction bias, i.e., even if no query is given, ICL still returns unbalanced results (Zhao et al., 2021) decoded from tokens of prompt template (see Appendix E for details).
@@
Shortcut Induction. Note that a $k$-shot ICL input sequence always contains a $(k-1)$-shot sequence, where the $k$-th forerunner (served as the $(k-1)$-shot query) is previously processed. So every forerunner can directly retrieve previously processed forerunners of demonstrations to copy their induction results directly. The non-trivial result (Table 2) with all forerunners disconnected from each other confirms such infer.

### v4

@@ 4 Induction Circuits in Large Language Models
This section mainly shows how LMs utilize the encoded linear text representations in induction circuits with a typical 2-step form (Singh et al., 2024b): Forerunner Token Heads merge the demonstration text representations in the forerunner tokens into their corresponding label tokens with a selectivity regarding the compatibility of demonstrations and label semantics. Induction Heads copy the information in the label representations similar to the query representation back to the query, which is conducted on task-specific subspaces, enabling LMs to solve multiple tasks by multiplexing hidden spaces.
@@ 4.1 Step 2, Forerunner Token Head: Copy from Text Feature to Label Token
This subsection mainly examines and measures the forerunner token heads, which copy the information in the forerunner token into label tokens. We focus on how the representations in these tokens are merged, especially when the semantics of demonstrations and corresponding labels are disjoint, to explain the robustness of ICL against wrong labels present in the demonstrations.
@@
Text Representations are Copied without Selectivity. For each attention head, we extract the attention score [^7] This paper use notation $\alpha_{K\rightarrow Q}$ to denote attention score with attention query $Q$ and attention key $K$, which is same with the flow of information. $\alpha_{s_{t}\rightarrow y_{t}}$ from the forerunner token $[s_{t}]$ (as attention key) to the corresponding label token $[y_{t}]$ (as attention query). We then mark the head with $\alpha_{s_{t}\rightarrow y_{t}}\geqslant 5/n_{t}$ ($n_{t}$: the length of tokens before $[y_{t}]$) as a Forerunner Token Head and count them in each layer. The results are shown in Fig. 5 (Middle, “Correct Label”), where the peak matches the copy period in Fig. 5 (Left). Moreover, to investigate the influence of the correctness of label tokens, we replace $[y_{t}]$ with a wrong label token [^8] For example, for a label space of “positive” and “negative”, if $[y_{t}]$ is “positive”, we replace it to “negative”., where the results in Fig. 5 (Middle, “Wrong Label”) are almost identical to the correct-label setting, suggesting that the forerunner token heads don’t show selectivity toward the semantic consistency between input text and labels, and simply merge the input text representations into the label tokens. Furthermore, we find (Appendix D) that the copy processing is inherent: During the copy processing, the preceding tokens are copied to the subsequent tokens, regardless of whether these copied tokens are forerunner tokens, which indicates that such copy processing is a universal inference behavior established during pre-training, rather than being evoked by the special tokens in the in-context learning (ICL) input, while still aiding ICL process (discussed in §5.1).
@@
Hidden States of Label Tokens are Joint Representations of Text Encodings and Label Semantics. Given the findings above, we probe the content of hidden states of label tokens $[y_{t}]$, i.e., how the copied text representation interacts with the original label semantics. We first train two centroid classifiers to predict the corresponding label $y_{t}$: (1) $\mathcal{C}_{s}$ trained on the hidden states of forerunner tokens $[s_{t}]$ and (2) $\mathcal{C}_{y}$ trained on the hidden states of label tokens $[y_{t}]$. To check whether the label tokens include the information of forerunner tokens, we use $\mathcal{C}_{s}$ to predict the label on the hidden state of label token $[y_{t}]$ in Fig. 5 (Right, solid). It shows that, during the copy processing, high classification accuracies can be achieved both on the correct label tokens and wrong label tokens, suggesting that the text features in the forerunner tokens can be partly and linearly detected in the label tokens. Moreover, results using $\mathcal{C}_{y}$ (dotted line) show extreme results, suggesting the label information remains in the label token. So, we can conclude that: Hidden states of label tokens are joint representations of label semantics and text representations. Moreover, interestingly, the accuracies from $\mathcal{C}_{s}$ shown in Fig. 5 (Right) decline after layer 35, suggesting that the information sharing between the forerunner tokens and the label tokens ends in later layers, which aligns with the results in Fig. 5 (Middle).
@@
Label Denoising is Conducted on the Overlap of Label Semantics and Text Representations. Notice that in Fig. 5 (Right, solid), compared to the typical results predicted on the forerunner tokens, accuracies are improved on the correct label and suppressed on the wrong label, which suggests that information consistent with the label semantics is easier to be enhanced by the label tokens and vice versa, showing a feature selectivity on the consistency between the text representations and the label semantics. Given the observation that the information on label semantics and text features can be extracted separately and linearly, we can confirm that these two kinds of information are located in different sub-spaces of the hidden states, and linearly merged by the attention operation of forerunner token heads. Moreover, given the fact that there is no selectivity is observed in the copy behavior of the forerunner token heads (Fig. 5 (Middle)), it is intuitive that the feature selectivity shown in Fig. 5 (Right) comes from the arithmetical interaction of feature vectors on the overlap of sub-spaces between the label semantics and text features, making ICL stable against label noise (Min et al., 2022). Moreover, as mentioned by Wei et al. (2023), large models show poorer stability against label noise, and we infer that the larger hidden dimensions in larger models lower the overlap between the sub-spaces of label semantics and text representations to reduce the interaction.
@@
Figure 5: Hidden states copy magnitude from forerunner tokens to label tokens against layers. Left: Kernel alignment between the forerunner token (the copy source) and the abstract label token of the next layer (the copy target). Middle: Curves: The count of marked forerunner token heads with correct and wrong labels; Colored Areas: The maximum attention scores from forerunner token to query (copy magnitude) with correct and wrong labels (detailed attention head statistical data is in Appendix H.1). Right: Centroid classifier results predicted on the hidden states of correct and wrong label tokens, on SST-2 and MR. Solid: Predicted by classifiers $\mathcal{C}_{s}$ trained on hidden states of forerunner tokens. Dotted: Predicted by classifiers $\mathcal{C}_{y}$ trained on hidden states of label tokens.
@@ 4.2 Step 3, Induction Head: Feature Retrieval on Task Subspace
This subsection examines the existence of the induction heads, which retrieve label token features similar to the queries’ forerunner token feature, and copy the retrieved features back to the query. We claim the necessity of multi-head attention in this process: Correct feature retrieval can only be conducted on the subspace of the hidden space, which is captured by some attention heads.
@@
Induction is Correct in Minority Subspaces. Similar to Fig. 5 (Middle), we mark (1) the attention heads with the sum of attention scores from all the label tokens in the demonstrations $[y_{1}],\dots,[y_{k}]$ (as attention keys) to the query’s forerunner token $[s_{q}]$ (as attention query) of more than $5k/n_{t}$ as Induction Heads, and (2) attention heads with the sum of scores from all the correct label tokens more than $5k/|\mathbb{Y}|n_{t}$ as Correct Induction Heads ($\mathbb{Y}$ is the label space). We show the number of both kinds of induction heads in Fig. 6 (Left, detailed head statistics in Appendix H.1), where an unimodal pattern is observed later than the copy processing of Step 2. Moreover, more than half of the induction heads are not correct ones, suggesting that task-specific feature similarity can only be caught on some induction subspaces (defined by low-rank transition matrix $W_{Q}^{h\top}W_{K}^{h}$ of correct induction head $h$). We enhance this claim in Fig. 6 (Middle) (details in Appendix A.4), where both vanilla attention (without transformation and head split) and attention scores averaged among all heads show low assignment on correct label tokens, while some heads show considerable correctness. Considering the average value, the majority of attention heads almost randomly copy label token information to the query, causing the final prediction biased to the frequency of labels present in the demonstrations (Zhao et al., 2021). As the reason, we infer that the hidden states are sufficient (Fig. 3) but not minimum for ICL, where redundant information interferes with the operation of induction heads.
@@
Figure 6: Measurements for induction heads. Left: The count of marked induction heads and correct induction heads against layers. Middle: The correctness of attention assignment (the sum of attention scores from query’s forerunner token towards correct label tokens normalized by attention scores towards all the label tokens) extracted from the following attention calculations: Vanilla Attention: attention scores directly calculated on full dimensionality. Head Average: the averaged attention scores among all the heads. Best Ind. Head: the scores of attention head with the most correctness. Right: The correct induction heads overlap of all dataset pairs.
@@
Figure 7: Label representations of $k$ demonstrations visualized on Left 4: correct and Right 4: wrong induction head, on one sample of SST-2 (see Appendix H.3). $\circ$: “positive” label, $\times$: “negative” label, $\blacktriangle$: zero vector. Color: Radial density of each categories, to estimate the attention assigned to query, negative to positive (cartography: Appendix A.5).
@@
Some Induction Subspaces are Task-specific. We check if different tasks share the same induction subspaces based on the overlap of the correct induction heads across different datasets. Given $n_{\mathcal{D}}(h)$ as the number of times $h$ is marked as correct induction head on dataset $\mathcal{D}$, the overlap rate $S$ is defined as:
$S(\mathcal{D}_{1},\mathcal{D}_{2})=\frac{2\sum_{\forall h}\mathrm{min}[n_{\mathcal{D}_{1}}(h),n_{\mathcal{D}_{2}}(h)]}{\sum_{\forall h}n_{\mathcal{D}_{1}}(h)+n_{\mathcal{D}_{2}}(h)}.$
@@
The results are shown in Fig. 6 (Right), where: (1) A significant overlap of induction heads indicates that a part of correct induction heads is inherent in the model, built by the pre-training process (Reddy, 2024; Singh et al., 2024b). (2) Such overlap is not fully observed, suggesting that some induction subspaces are task-specific: Input texts evoke task-specific induction heads, enabling the anisotropy multiplex of different subspaces in the hidden spaces to transmit relevant information for various tasks. Therefore, we can analogize ICL as an implicit end-to-end multi-task learning with hidden state multiplexing since multi-task learning also utilizes various task heads on common bottom network layers and informative hidden states (Zhang & Yang, 2021).
@@
Demonstrations Saturate on Induction Subspace. We visualize the demonstrations’ label token representations mapped on the induction subspaces by the transition matrix $W_{Q}^{h\top}W_{K}^{h}$ and principal component analysis in Fig. 7, indicating: (1) Compared to correct induction heads (Left 4), wrong induction heads (Right 4) are easier to map label representations linear-inseparably. (2) In the early stage of demonstration feeding ($k=1\rightarrow 2$), when a new demonstration is given, the morphology of attention assignment towards query updates significantly (shown as background color in Fig. 7, estimated by the radial density of various labels; see §A.5 for details), while in the late stage ($k=15\rightarrow 16$), attention assignment morphology is stable. This can explain the demonstration saturation (Agarwal et al., 2024; Bertsch et al., 2024): ICL performance is submodular against the number of demonstrations. Intuitively, since demonstrations follow a prior distribution, representation of a new demonstration is likely to be located within the closure of existing demonstrations, making it less diverse to contribute to the attention assignment in the induction subspace.
@@ 5 Label $y_{i}$rightarrow Query Forerunner $s_{q}$
To (1) examine the causality between the attention connection specified by the 3 steps and ICL inference, and (2) demonstrate that our 3-phase circuit dominates or at least participates in ICL process, we disconnect the related attention connection of each step in the proposed circuit (see Appendix A.7 for details), and test the accuracies without such connections, as shown in Table 1. The results indicate that, compared to the controlled results, where trivial connections are removed, the accuracy of ICL significantly decreases when the non-trivial connections designated by the proposed circuit are ablated. This supports the existence of our circuit. However, the result doesn’t fully match expectations, for example, the result without induction (line 5) should be consistent with the zero-shot inference result (line 6), since all the expected communication from demonstration to query is intercepted, but the real result is better; and the contribution of Step 1 in later layers are unexpectedly high, indicating the existence of some bypass mechanisms parallelly contributing to ICL accuracies.
@@ 5.2 Bypass Mechanism
Parallel Circuits. Multiple 3-step circuits can execute in parallel, and one layer can assign multiple inference functions to different heads, causing dispersion and deserialization as shown in Fig. 8, where a narrow (fewer heads per layer) and deep (more layers) model is more likely to generate localized inference, and vice versa.
Direct Decoding. Residual connection to the LM heads allows intermediate hidden states to be decoded directly. Intuitively, a shortcut from Step 1 encodings to the LM head enables ICL with zero-shot capacities, since we have confirmed that encoded representations are informative for ICL tasks (§3.2), while the decoding methods should be selected carefully (Cho et al., 2024) (e.g. with essential calibration). On the other hand, shortcuts from insufficiently encoded features may lead to meaningless information decoded by language model heads, causing prediction bias, i.e., even if no query is given, ICL still returns unbalanced results (Zhao et al., 2021) decoded from tokens of prompt template (see Appendix E for details).
@@
Shortcut Induction. Note that a $k$-shot ICL input sequence always contains a $(k-1)$-shot sequence, where the $k$-th forerunner token (served as the query of $(k-1)$-shot sequence) is previously processed by the $(k-1)$-shot inference. So every forerunner can directly retrieve previously processed forerunner tokens in the demonstrations to copy their induction results directly. The non-trivial result (Table 2) with all forerunner tokens disconnected from each other confirms such infer.